# Wi-Fi Radar via Over-the-Air Referencing: Bridging Wi-Fi Sensing and Bistatic Radar

### v2

@@ I Introduction
Early Wi-Fi sensing studies primarily relied on coarse metrics such as received signal strength indicator (RSSI) [5, 6]. A major turning point was the release of the Linux 802.11n channel state information (CSI) tool [7], which enabled researchers to access CSI from commercial off-the-shelf (COTS) IEEE 802.11n devices. In this context, CSI refers to the complex-valued channel frequency response (CFR) measured across orthogonal frequency division multiplexing (OFDM) subcarriers. From a communication-system perspective, CSI is obtained for communication purposes such as equalization, precoding, and link adaptation, rather than being designed specifically for sensing purposes. Throughout this paper, the terms CSI and CFR are used interchangeably: CSI is used when referring to channel measurement and acquisition, while CFR denotes the frequency-domain representation of the channel and its Fourier-transform relationship to the channel impulse response (CIR), which represents the channel in the delay domain.
This CSI acquisition capability has triggered extensive research on CSI-based Wi-Fi sensing, including motion-related sensing based on temporal variations of the wireless channel, as summarized in representative surveys such as [2, 3, 4]. More recently, Wi-Fi sensing has matured beyond ad-hoc research prototypes, culminating in the standardization of IEEE 802.11bf [8], which defines channel measurement procedures to explicitly support a wide range of sensing use cases, including range, motion, and gesture detection, within Wi-Fi networks.
@@
Alongside the aforementioned Wi-Fi sensing approaches, there also exists a line of research that exploits COTS Wi-Fi transmissions for radar applications [9, 10]. For clarity, we refer to methods operating without explicit reference signals as Wi-Fi sensing, and those employing explicit reference signals for phase-coherent analysis as Wi-Fi radar. Early Wi-Fi radar studies typically relied on measurement-grade receivers and external reference acquisition mechanisms, such as reference antennas or wired signal taps, as illustrated in Fig. 1(a) and (b), to enable coherent delay or Doppler estimation [9, 11, 10, 12, 13]. More recent works have demonstrated that reference signals can also be exploited in fully COTS-based systems to stabilize CSI measurements, i.e., the phase of the underlying CFR [14, 15, 16, 17]. These reference-aided architectures are fundamentally consistent with bistatic radar principles and clearly demonstrate the potential of Wi-Fi signals for radar-like sensing. However, due to their reliance on cable routing or dedicated reference antenna placement, most prior Wi-Fi radar functionalities cannot be integrated into typically deployed Wi-Fi systems.
These observations reveal a fundamental gap between existing Wi-Fi sensing and Wi-Fi radar approaches. While reference-aided Wi-Fi radar enables physically meaningful delay and Doppler analysis, such architectures cannot be integrated into typical Wi-Fi systems. Conversely, Wi-Fi sensing approaches that rely, for example, on CFR magnitude alone do not require any modification to existing Wi-Fi systems, but lack direct access to physically interpretable propagation parameters.
@@
This paper is organized as follows. Section II reviews related work on Doppler-based Wi-Fi sensing and reference-aided Wi-Fi radar. Section III discusses the reference ambiguity due to independent clocks in practical Wi-Fi systems. Section IV presents the proposed phase-coherent CIR construction framework, including delay and phase referencing based on a stable LoS-path reference. Section V demonstrates bistatic range estimation through human gait and respiration experiments. Section VI presents delay–Doppler response estimation results for moving targets. Section VII provides additional discussion, followed by Section VIII, which concludes the paper.
@@ II Related Work
We review prior Wi-Fi sensing studies relevant to this work, followed by reference-aided Wi-Fi radar approaches. Estimation of propagation distance or time-of-flight based on statistical or geometry-based features of CSI [19, 20] is outside the scope of this review, as such studies primarily aim at inferring static distance or position from channel statistics, rather than resolving delay–Doppler responses from phase-coherent CIRs.
@@ II-A Doppler-Based Wi-Fi Sensing
In [23], respiration sensing is performed by selecting subcarriers whose CFR magnitudes exhibit strong target-induced periodic variations. Spectral analysis is then applied to the selected magnitude-domain time series to estimate the respiration rate.
WiFiU [24] adopts a principal component analysis (PCA)-based dimensionality reduction of the CFR magnitude, followed by STFT on the resulting PC scores to extract Doppler-related information. Widar [25] further interprets the dominant spectral components as path length change rates, and estimates target velocity and location through an explicit geometric model. It is worth noting that, in both approaches, the extracted Doppler-related quantities are derived from magnitude-domain spectral analysis.
In particular, Li et al. [26] provide an insightful analysis closely related to the lack of stable phase reference addressed in this work. The authors contrast Wi-Fi sensing and Wi-Fi radar by clarifying what physical quantities are actually estimated in each case. They show that magnitude-based CFR processing inherently loses Doppler sign information and physically meaningful phase evolution because magnitude-only processing removes the phase evolution induced by propagation path-length changes. As a result, the extracted spectral components reflect periodic signal strength fluctuations rather than a true Doppler frequency defined by wave propagation.
@@ II-B Reference-Aided Wi-Fi Radar
Table I summarizes the positioning of this work with respect to representative Wi-Fi sensing and reference-aided Wi-Fi radar studies, highlighting differences in Tx and Rx types, reference signals, ISAC-enabling capability, and supported standards. Here, ISAC-enabling indicates whether such functionality can be realized within typically deployed Wi-Fi systems without additional reference hardware. In [9, 11], a bistatic radar architecture is adopted, where a COTS Wi-Fi Tx and a measurement-grade Rx are used. The Rx employs separate reference and surveillance antennas, as illustrated in Fig. 1(b), with the antenna configuration designed such that target-induced variations are primarily captured in the Tx–surveillance path. The target-induced Doppler frequency is then extracted by correlating the reference and surveillance signals.
As an alternative reference acquisition strategy, [10, 12, 13] inserted a directional coupler between a COTS Wi-Fi Tx and its antenna to obtain a wired reference signal, as illustrated in Fig. 1(a). It should be noted that in all these studies [9, 11, 10, 12, 13], the Rx is instrument-grade, and sensing is performed directly on received signals rather than on CSI measurements obtained from COTS Wi-Fi devices. Moreover, these studies largely predate the widespread availability of CSI on COTS Wi-Fi hardware, which emerged later with IEEE 802.11n and CSI extraction tools [7]. In this sense, Wi-Fi is exploited merely as a signal of opportunity, without leveraging the CSI natively observed by COTS Wi-Fi devices. Nevertheless, these works clearly demonstrate that Wi-Fi signals can fundamentally support radar functionality.
@@ III CIR Representation and Reference Ambiguity Due to Free-Running Clocks
This section briefly reviews the CIR representation and highlights the reference ambiguity arising from free-running clocks in practical Wi-Fi systems, which fundamentally prevents phase-coherent analysis and physically meaningful time-domain processing.
@@ III-A CIR Representation and Doubly Selective Channels
The equivalent low-pass representation of the CIR can be described using a tapped delay line (TDL) model [27]. Under this model, the CIR at a given measurement time $t$ is expressed as a superposition of discrete propagation paths, each characterized by a complex path gain and a propagation delay, as
@@
where $L$ denotes the number of discrete propagation paths, $\eta_{\ell}$ and $\tau_{\ell}$ represent the complex amplitude and delay of the $\ell$-th path, respectively, and $f_{\mathrm{c}}$ is the carrier frequency. Here, $\tau$ denotes the signal propagation delay measured from the transmission time, and $t$ denotes the observation time of the CIR, corresponding to successive channel measurements rather than a continuous-time variable. This representation captures the multipath structure of the wireless channel in the delay domain, where the complex phase of each path directly reflects the propagation path length.
In dynamic environments, the underlying wireless channel is doubly selective, exhibiting variations across both delay and time. The CIR $h(\tau,t)$ provides a delay–time representation that captures such joint delay–time variations.
@@ III-B Reference Ambiguity Due to Free-Running Clocks
In Wi-Fi systems, where the Tx and Rx employ independent, free-running clocks, neither a common timing reference nor an absolute phase reference is available between them. As a result, the CIR observed at the Rx does not preserve the absolute delay reference to the transmission timing, nor does it retain a phase-coherent representation across successive observations. Accordingly, the observed CIR $h_{\mathrm{raw}}(\tau,t)$ is related to the underlying physical CIR $h(\tau,t)$ through an unknown delay offset and an unknown phase offset, and is given by
@@
where $\alpha_{t}$ and $\beta_{t}$ denote the delay offset and phase offset, respectively, which may vary across observation times $t$ [21, 22]. These offsets originate from free-running clocks at the Tx and Rx, and prevent direct use of absolute delay and phase information, thereby precluding phase-coherent analysis of the doubly selective channel in practical Wi-Fi systems without a shared reference. Note that the delay offset $\alpha_{t}$ of the devices used in this study has been experimentally characterized in [17].
When (2) is viewed in the frequency domain, the resulting complex-valued CFR $H_{\mathrm{raw}}(f,t)$, where $f$ denotes the OFDM subcarrier frequency, is inherently discontinuous with respect to time $t$ due to these unknown offsets, which precludes direct temporal or spectral analysis. To circumvent this issue, many Wi-Fi sensing studies, including [23], operate on the magnitude of the CFR, $|H_{\mathrm{raw}}(f,t)|$. Taking the magnitude effectively removes these unknown offsets, yielding an apparently smooth time series that is amenable to spectral analysis. This convenience, however, comes at the cost of discarding the phase evolution associated with wave propagation, thereby precluding any phase-coherent interpretation of propagation delay or Doppler frequency.
@@
Figure 2: Processing flow from measured CFR to the delay–Doppler response $s(\tau,\nu;t)$. The LoSRef exploits the Tx–Rx line-of-sight path as a physical delay and phase reference, enabling phase-coherent alignment of CIR across packet acquisitions.
@@ IV Phase-Coherent Time-Variant Impulse Response Estimation
The offsets $\alpha_{t}$ and $\beta_{t}$ in (2) are unknown and vary across observation times, which prevents phase-coherent alignment of CIRs across successive packet acquisitions. In contrast, the delay of the LoS-path component in (1) is determined solely by the Tx–Rx distance and is therefore expected to remain constant as long as the Tx and Rx are stationary. Exploiting this temporal stability, the LoSRef uses the LoS-path component as an OTA reference to mitigate the effects of these unknown offsets and to enable phase-coherent time-variant CIR estimation. Fig. 2 summarizes the overall processing flow of the proposed framework, incorporating LoSRef, from the measured CFR to the delay–Doppler response, and each processing block is detailed in the following subsections.
@@ IV-A CIR Construction for Robust LoS-Path Identification
It should be noted that the resulting CIR includes the phase and gain of the transceivers and the radio channel, and is convolved with the impulse response of the applied frequency-domain window. For simplicity, this windowed version is hereafter referred to as the CIR. Similarly, although the observations are discrete and the transformation should formally be written as an inverse discrete Fourier transform (IDFT), we retain the simplified representation in (3) to avoid unnecessary notational complexity.
To construct a stable delay-domain CIR, frequency-domain windowing and delay-domain oversampling are applied, following the same principle as in our prior Wi-Fi channel sounding work [17, Section III-C], with window and oversampling parameters adapted to the present setup. Specifically, a Blackman window is used prior to the inverse Fourier transform to suppress delay-domain sidelobes. In addition, delay-domain oversampling with a factor of $\kappa=32$ is performed via zero padding in the frequency domain. It is emphasized that such oversampling does not improve the intrinsic delay resolution, which remains fundamentally limited by the signal bandwidth ($1/B=$6.3\text{\,}\mathrm{ns}$\$), but is nevertheless essential for achieving accurate and temporally stable peak detection under discretized delay sampling.
@@ IV-B LoSRef-Based Delay Calibration and Phase Alignment
Let the known LoS OTA Tx–Rx propagation distance $d_{\text{Tx--Rx}}$ be used as the reference distance $d_{\mathrm{ref}}$. For clarity, this OTA LoS-path referencing scheme is hereafter simply referred to as LoSRef. While the measurement accuracy of $d_{\mathrm{ref}}$ may warrant careful discussion depending on the target application, this issue is beyond the scope of this work, since $d_{\mathrm{ref}}$ is used solely for absolute delay-axis alignment. Doppler frequency estimation, which is based on the temporal derivative of the propagation phase, is therefore independent of the absolute value of $d_{\mathrm{ref}}$, as long as it is constant over time.
The propagation delay of the LoS path relative to the transmitted signal is given by $\tau_{\mathrm{ref}}=d_{\mathrm{ref}}/c$, where $c$ denotes the speed of light. In an ideal delay-synchronized system, the LoS-path component would therefore appear at $\tau_{\mathrm{ref}}$ in the CIR.
In practice, however, this alignment is not observed, because the Tx and Rx are not delay-synchronized and the delay axis of the measured CIR is not referenced to the actual packet transmission timing. As a result, the absolute delay origin varies across packet acquisitions.
To address this issue, the delay axis of the CIR is calibrated by identifying the dominant component, assumed to correspond to the LoS path, and aligning its delay with the reference delay $\tau_{\mathrm{ref}}$. This procedure establishes a common delay reference across packet acquisitions. [^1] Strictly speaking, the identified peak corresponds to a delay bin dominated by the LoS path rather than an isolated propagation path. Specifically, the delay calibration is performed to satisfy
@@
Figure 3: LoSRef: OTA LoS-path referencing for delay calibration and phase alignment. The dominant LoS-path component is used as an OTA reference to compensate unknown offsets arising from free-running clocks at the Tx and Rx.
@@
as illustrated in Fig. 3. The above calibration procedure follows the same mathematical formulation as our prior channel sounding work [17].
In addition, the magnitude of the CIR is normalized such that $\lvert h(\tau_{\mathrm{ref}},t)\rvert$ matches the free-space path gain corresponding to the reference distance $d_{\mathrm{ref}}$. Consequently, the resulting CIR $h(\tau,t)$ represents the complex path gain associated with each propagation path.
In this calibration procedure, the delay shift by $\tau_{\mathrm{ref}}-\tau_{t}^{\star}$ is referred to as delay calibration, while the phase rotation by ${\rm e}^{-{\rm j}\theta_{t}^{\star}}$ is referred to as phase alignment. Through phase alignment, the phase of each multipath component is expressed relative to that of the dominant (LoS-path) component. As a result, the phase of the CIR becomes continuous across successive acquisitions, enabling phase-coherent time-variant and Doppler analysis.

### v3

@@ I Introduction
Early Wi-Fi sensing studies relied on coarse metrics such as received signal strength indicator (RSSI) [5, 6]. A major turning point was the release of the Linux 802.11n channel state information (CSI) tool [7], which enabled researchers to access CSI from commercial off-the-shelf (COTS) IEEE 802.11n devices. In this context, CSI refers to the complex-valued channel frequency response (CFR) measured across orthogonal frequency division multiplexing (OFDM) subcarriers. From a communication-system perspective, CSI is obtained for communication purposes such as equalization, precoding, and link adaptation, rather than being designed for sensing purposes. Throughout this paper, the terms CSI and CFR are used interchangeably: CSI is used when referring to channel measurement, while CFR denotes the frequency-domain representation of the channel and its Fourier-transform relationship to the channel impulse response (CIR), which represents the channel in the delay domain.
This CSI acquisition capability has triggered research on CSI-based Wi-Fi sensing, including motion-related sensing based on temporal variations of the wireless channel, as summarized in surveys such as [2, 3, 4]. More recently, Wi-Fi sensing has matured, culminating in the standardization of IEEE 802.11bf [8], which defines channel measurement procedures to support sensing use cases, including range, motion, and gesture detection.
@@
Alongside the Wi-Fi sensing approaches, there exists a line of research that exploits COTS Wi-Fi transmissions for radar applications [9, 10]. For clarity, we refer to methods operating without explicit reference signals as Wi-Fi sensing, and those employing explicit reference signals for phase-coherent analysis as Wi-Fi radar. Early Wi-Fi radar studies typically relied on measurement-grade receivers and external reference acquisition mechanisms, such as reference antennas or wired signal taps, as illustrated in Fig. 1(a) and (b), to enable coherent delay or Doppler estimation [9, 11, 10, 12, 13]. Recent works have demonstrated that reference signals can also be exploited in fully COTS-based systems to stabilize CSI measurements, i.e., the phase of the CFR [14, 15, 16, 17]. These reference-aided architectures are consistent with bistatic radar principles and demonstrate the potential of Wi-Fi signals for radar-like sensing. However, due to their reliance on cable routing or dedicated reference antenna placement, most prior Wi-Fi radar functionalities cannot be integrated into typically deployed Wi-Fi systems.
These observations reveal a fundamental gap between Wi-Fi sensing and Wi-Fi radar approaches. While reference-aided Wi-Fi radar enables physically meaningful delay and Doppler analysis, such architectures cannot be integrated into typical Wi-Fi systems. Conversely, Wi-Fi sensing approaches that rely, for example, on CFR magnitude alone do not require any modification to existing Wi-Fi systems, but lack direct access to physically interpretable propagation parameters.
@@
This paper is organized as follows. Section II reviews related work on Doppler-based Wi-Fi sensing and reference-aided Wi-Fi radar. Section III discusses the reference ambiguity due to independent clocks in practical Wi-Fi systems. Section IV presents the proposed phase-coherent CIR construction framework, including delay and phase referencing based on a stable LoS-path reference. Section V demonstrates bistatic range estimation through human gait and respiration experiments. Section VI presents delay–Doppler response estimation results for moving targets. Section VII provides concluding remarks.
@@ II Related Work
We review Wi-Fi sensing studies relevant to this work, followed by reference-aided Wi-Fi radar approaches. Estimation of propagation distance or time-of-flight based on statistical or geometry-based features of CSI [19, 20] is outside the scope of this review, as such studies primarily aim at inferring static distance or position from channel statistics, rather than resolving delay–Doppler responses from phase-coherent CIRs.
@@ II-A Doppler-Based Wi-Fi Sensing
In [23], respiration sensing is performed by selecting subcarriers whose CFR magnitudes exhibit target-induced periodic variations. Spectral analysis is then applied to the selected magnitude-domain time series to estimate the respiration rate.
WiFiU [24] adopts a principal component analysis (PCA)-based dimensionality reduction of the CFR magnitude, followed by STFT on the PC scores to extract Doppler-related information. Widar [25] further interprets the dominant spectral components as path length change rates, and estimates target velocity and location through an explicit geometric model. In both approaches, the Doppler-related quantities are derived from magnitude-domain spectral analysis.
In particular, Li et al. [26] provide an insightful analysis closely related to the lack of stable phase reference addressed in this work. The authors contrast Wi-Fi sensing and Wi-Fi radar by clarifying what physical quantities are estimated in each case. They show that magnitude-based CFR processing inherently loses Doppler sign information and physically meaningful phase evolution because magnitude-only processing removes the phase evolution induced by propagation path-length changes. As a result, the extracted spectral components reflect periodic signal strength fluctuations rather than a true Doppler frequency defined by wave propagation.
@@ II-B Reference-Aided Wi-Fi Radar
Table I summarizes the positioning of this work with respect to representative Wi-Fi sensing and reference-aided Wi-Fi radar studies, highlighting differences in Tx and Rx types, reference signals, ISAC-enabling capability, and supported standards. Here, ISAC-enabling indicates whether such functionality can be realized within typically deployed Wi-Fi systems without additional reference hardware. In [9, 11], a bistatic radar architecture is adopted, where a COTS Wi-Fi Tx and a measurement-grade Rx are used. The Rx employs separate reference and surveillance antennas, as illustrated in Fig. 1(b), with the antenna configuration such that target-induced variations are captured in the Tx–surveillance path. The target-induced Doppler frequency is then extracted by correlating the reference and surveillance signals.
As an alternative reference acquisition strategy, [10, 12, 13] inserted a directional coupler between a COTS Wi-Fi Tx and its antenna to obtain a wired reference signal, as illustrated in Fig. 1(a). It should be noted that in all these studies [9, 11, 10, 12, 13], the Rx is instrument-grade, and sensing is performed directly on received signals rather than on CSI measurements obtained from COTS Wi-Fi devices. Moreover, these studies largely predate the widespread availability of CSI on COTS Wi-Fi hardware, which emerged later with IEEE 802.11n and CSI extraction tools [7]. In this sense, Wi-Fi is exploited merely as a signal of opportunity, without leveraging the CSI natively observed by COTS Wi-Fi devices. Nevertheless, these works demonstrate that Wi-Fi signals can fundamentally support radar functionality.
@@ III CIR Representation and Reference Ambiguity Due to Free-Running Clocks
This section discusses the reference ambiguity in practical Wi-Fi systems, which fundamentally prevents phase-coherent analysis and physically meaningful time-domain processing.
@@ III-A CIR Representation and Doubly Selective Channels
The equivalent low-pass representation of the CIR can be described using a tapped delay line (TDL) model [27]. Under this model, the CIR at measurement time $t$ is expressed as a superposition of discrete propagation paths, each characterized by a complex path gain and a propagation delay, as
@@
where $L$ denotes the number of paths, $\eta_{\ell}$ and $\tau_{\ell}$ represent the complex amplitude and delay of the $\ell$-th path, respectively, and $f_{\mathrm{c}}$ is the carrier frequency. Here, $\tau$ denotes the signal propagation delay measured from the transmission time, and $t$ denotes the observation time of the CIR, corresponding to successive channel measurements rather than a continuous-time variable. This representation captures the multipath structure of the wireless channel in the delay domain, where the phase of each path reflects the propagation path length.
In dynamic environments, the underlying wireless channel is doubly selective, exhibiting variations across both delay and time. The CIR $h(\tau,t)$ captures such joint delay–time variations.
@@ III-B Reference Ambiguity Due to Free-Running Clocks
In Wi-Fi systems, where the Tx and Rx employ independent, free-running clocks, neither a common timing reference nor an absolute phase reference is available between them. As a result, the CIR observed at the Rx does not preserve the absolute delay reference to the transmission timing, nor does it retain a phase-coherent representation across successive observations. Accordingly, the observed CIR $h_{\mathrm{raw}}(\tau,t)$ is related to the physical CIR $h(\tau,t)$ through an unknown delay offset and an unknown phase offset, and is given by
@@
where $\alpha_{t}$ and $\beta_{t}$ are the delay and phase offsets, respectively, which may vary across observation times $t$ [21, 22]. These offsets originate from free-running clocks at the Tx and Rx, and prevent direct use of absolute delay and phase information, thereby precluding phase-coherent analysis of the doubly selective channel in practical Wi-Fi systems without a shared reference. Note that the delay offset $\alpha_{t}$ of the devices used in this study has been experimentally characterized in [17].
When (2) is viewed in the frequency domain, the resulting complex-valued CFR $H_{\mathrm{raw}}(f,t)$, where $f$ denotes the OFDM subcarrier frequency, is inherently discontinuous with respect to time $t$ due to these unknown offsets, which precludes direct temporal or spectral analysis. To circumvent this issue, many Wi-Fi sensing studies, including [23], operate on the magnitude of the CFR, $|H_{\mathrm{raw}}(f,t)|$. Taking the magnitude removes these unknown offsets, yielding a smooth time series that is amenable to spectral analysis. This convenience, however, comes at the cost of discarding the phase evolution associated with wave propagation, thereby precluding any phase-coherent interpretation of propagation delay or Doppler frequency.
@@
Figure 2: Processing flow from measured CFR to the delay–Doppler response. The LoSRef exploits the Tx–Rx LoS path as a delay and phase reference, enabling phase-coherent alignment of CIR across acquisitions.
@@ IV Phase-Coherent Time-Variant Impulse Response Estimation
The offsets $\alpha_{t}$ and $\beta_{t}$ in (2) are unknown and vary across observation times, which prevents phase-coherent alignment of CIRs across successive packet acquisitions. In contrast, the delay of the LoS-path component in (1) is determined by the Tx–Rx distance and is therefore expected to remain constant as long as the Tx and Rx are stationary. Exploiting this temporal stability, the LoSRef uses the LoS-path component as an OTA reference to mitigate the effects of these unknown offsets and to enable phase-coherent time-variant CIR estimation. Fig. 2 summarizes the processing flow of the proposed framework, incorporating LoSRef, from the measured CFR to the delay–Doppler response, and each processing block is detailed in the following subsections.
@@ IV-A CIR Construction for Robust LoS-Path Identification
The resulting CIR includes the phase and gain of the transceivers and the radio channel, and is convolved with the impulse response of the applied frequency-domain window. For simplicity, this windowed version is hereafter referred to as the CIR. Similarly, although the observations are discrete and the transformation should be written as an inverse discrete Fourier transform (IDFT), we retain the simplified representation in (3) to avoid notational complexity.
To construct a stable delay-domain CIR, frequency-domain windowing and delay-domain oversampling are applied, following the same principle as in our prior Wi-Fi channel sounding work [17], with window and oversampling parameters adapted to the present setup. Specifically, a Blackman window is used prior to the inverse Fourier transform to suppress delay-domain sidelobes. In addition, delay-domain oversampling with a factor of $\kappa=32$ is performed via zero padding in the frequency domain. It is emphasized that such oversampling does not improve the intrinsic delay resolution, which remains limited by the signal bandwidth ($1/B=$6.3\text{\,}\mathrm{ns}$\$), but is nevertheless essential for achieving accurate and temporally stable peak detection under discretized delay sampling.
@@ IV-B LoSRef-Based Delay Calibration and Phase Alignment
Let the known LoS OTA Tx–Rx propagation distance $d_{\text{Tx--Rx}}$ be used as the reference distance $d_{\mathrm{ref}}$. For clarity, this OTA LoS-path referencing scheme is hereafter simply referred to as LoSRef. While the measurement accuracy of $d_{\mathrm{ref}}$ may warrant careful discussion depending on the target application, this issue is beyond the scope of this work, since $d_{\mathrm{ref}}$ is used solely for absolute delay-axis alignment. Doppler frequency estimation, which is based on the temporal derivative of the propagation phase, is therefore independent of the absolute value of $d_{\mathrm{ref}}$, as long as it is constant over time. The propagation delay of the LoS path relative to the transmitted signal is given by $\tau_{\mathrm{ref}}=d_{\mathrm{ref}}/c$, where $c$ denotes the speed of light. In an ideal delay-synchronized system, the LoS-path component would therefore appear at $\tau_{\mathrm{ref}}$ in the CIR.
In practice, however, this alignment is not observed, because the Tx and Rx are not delay-synchronized and the delay axis of the measured CIR is not referenced to the actual packet transmission timing. As a result, the absolute delay origin varies across packet acquisitions. To address this issue, the delay axis of the CIR is calibrated by identifying the dominant component, assumed to correspond to the LoS path, and aligning its delay with the reference delay $\tau_{\mathrm{ref}}$. This procedure establishes a common delay reference across packet acquisitions. [^1] Strictly speaking, the identified peak corresponds to a delay bin dominated by the LoS path rather than an isolated propagation path. Specifically, the delay calibration is performed to satisfy
@@
Figure 3: LoSRef: OTA LoS-path referencing for delay calibration and phase alignment. The dominant LoS-path component is used as an OTA reference to compensate unknown offsets arising from free-running clocks.
@@
as illustrated in Fig. 3. The above calibration procedure follows the same mathematical formulation as our prior channel sounding work [17]. In addition, the magnitude of the CIR is normalized such that $\lvert h(\tau_{\mathrm{ref}},t)\rvert$ matches the free-space path gain corresponding to the reference distance $d_{\mathrm{ref}}$. The resulting CIR $h(\tau,t)$ represents the complex path gain associated with each propagation path. In this calibration procedure, the delay shift by $\tau_{\mathrm{ref}}-\tau_{t}^{\star}$ is referred to as delay calibration, while the phase rotation by ${\rm e}^{-{\rm j}\theta_{t}^{\star}}$ is referred to as phase alignment. Through phase alignment, the phase of each multipath component is expressed relative to that of the dominant (LoS-path) component. As a result, the phase of the CIR becomes continuous across successive acquisitions, enabling phase-coherent time-variant and Doppler analysis.